4. Suppose you have a sample of CO2 in gas-tight syringe. The gas volume is 25.0 mL at 20.0°C. What is the final volume of the gas if you hold the syringe in your hand to raise the temperature to 37°C? (26.5 mL)


1
Expert's answer
2021-04-19T05:10:52-0400

"\frac{V_1}{T_1} = \frac{V_2}{T_2}"


V1 = 25 mL , T1 = 20⁰C , V2 = ? , T2 = 37⁰C


"\frac{25mL}{20⁰C} = \frac{V_2}{37⁰C}"


V2 = 46.25 mL is the final volume